OpenStudy (anonymous):

Math Analysis: Find all values of x between 0 degrees and 360 degrees that satisfy the equation. Round the approximate answers to the nearest tenth of a degree. 3. 3cot^2x -1 = 0 4. 2cos^2x + 9cosx + 4 = 0

OpenStudy (anonymous):

so i use the formula for tan2x?

OpenStudy (anonymous):

it's cot^2 x

OpenStudy (anonymous):

\[\cot ^{2} x \]

OpenStudy (kropot72):

Good. Sorry for my error. Lets start again.

OpenStudy (anonymous):

\[3\cot ^{2}x - 1 = 0\]

OpenStudy (anonymous):

\[2\cos ^{2}x + 9\cos + 4 = 0\]

OpenStudy (kropot72):

Factorising: \[3\cot ^{2}x-1=(\sqrt{3}\cot x+1)(\sqrt{3}\cot x-1)\]=0 \[\sqrt{3}\cot x=1\] \[\sqrt{3}\cot x=-1\] Can you find x for both roots ? Hint:\[\cot x=\frac{1}{\tan x}\]

OpenStudy (anonymous):

why did you put a square root on the 3?

OpenStudy (kropot72):